# Fostering Collective Discourse: A Distributed Role-Based Approach to Online News Commenting

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: X››XXXXXX.XXXXXXXconference: In submission to CHI 2026; June 03–05, 2018; Woodstock, NYccs: Human-centered computing Collaborative and social computing systems and tools
@@ 1. Introduction
Previous studies have explored how the design of commenting systems can shape user behavior (Kiskola et al., 2021; Seering et al., 2019a). In many current news outlets, standardized commenting systems are used, where users simply post their comments or reply to others with limited opportunities to interact with other participants. This design often results in comments that are isolated expressions of personal thoughts, making it difficult for meaningful discussions to develop. As a result, discussions tend to unfold as a series of individual expressions, with users primarily focused on sharing their own views rather than collaborating toward a collective understanding or outcome. This behavior leads to fragmented conversations (Dahlberg, 2007b), where the overall discussion lacks depth. Consequently, similar viewpoints are often amplified, while diverse perspectives remain underrepresented (Anderson et al., 2014).
@@
Our research contributes to the exploration of an untapped space in designing collective online discourse by introducing a collaborative approach centered on “distributed roles,” reimagining how commenters can collectively shape and moderate discussions. We design and implement a novel system grounded in insights from previous works. Additionally, through experimentation, we present detailed findings on how our system differently shapes user behavior, understanding and identifying the newly emerged behavioral patterns using our suggested system. Finally, through our design exploration, we identify key design considerations and trade-offs associated with the values embedded in our system, which can serve as guidance for future efforts in designing collective discourse systems.
@@ 2. Related Work
By inviting reader comments, today’s news media have embraced greater user involvement in the journalistic process. This phenomenon has been explored by scholars through frameworks such as “participatory journalism” (Singer et al., 2011) and “reciprocal journalism” (Lewis et al., 2014); the audience has also become an active contributor to content. While there were positive views of news commenting acting as the digital cafés of a Public Sphere 2.0 (Ruiz et al., 2011), journalists and news outlets were often concerned about quality control, manageability, and the maintenance costs of user-generated content (Diakopoulos and Naaman, 2011).
@@ 2.2. Improving Discussion Quality through Moderation
A pre-moderation approach seeks to improve the discussion quality by addressing potential issues before they arise. These proactive approaches range from using design elements to nudge users toward prosocial participation (Seering et al., 2019a), to more direct interventions that review and approve comments before they appear publicly. Common design elements include explicitly displaying community rules and guidelines (Kraut, 2012; Matias, 2019), using prompts to encourage more thoughtful participation (McInnis et al., 2016), and employing curation strategies that influence users’ deliberation (McInnis et al., 2018). More direct strategies involve using algorithmic support to provide risk information to users (Chang et al., 2022) or moderators (Schluger et al., 2022), helping in the identification and resolution of tensions before they escalate.
The post-moderation approach, by contrast, takes place after content has been published. It focuses on identifying and mitigating the impact of harmful or inappropriate material through reactive review of comments. While traditional methods involve human moderators reviewing flagged content (Roberts, 2019; Seering et al., 2019b), the scale and volume of content necessitates heavy reliance on community participation to flag or report inappropriate comments.  Community-driven methods, also called a crowd-based approach, have proven to be effective for filtering and evaluating the quality of comments (Diakopoulos and Naaman, 2011). This approach leverages the collective judgment of the community to assess and organize comments, fostering a more dynamic and responsive environment. Collaborative platforms such as Wikipedia exemplify this, where community members work together to resolve disputes (Im et al., 2018) and contribute to the creation of high-quality articles (Zhang et al., 2017a). Another notable example is the moderation system used by Slashdot (Lampe et al., 2007), where users rate and moderate comments, with these decisions aggregated to determine comment quality. Slashdot’s effectiveness lies in its ability to harness the collective intelligence of its users, allowing for real-time participation in organizing and structuring discussions. This distributed form of moderation is both scalable and adaptable, addressing the diverse needs of its community.
@@ 2.3. Restructuring and Reimagining the Discussion Space
The deliberative model of democracy emphasizes the importance of reasoned argumentation, mutual respect, and the willingness to consider others’ perspectives as essential components for effective discourse (Dahlberg, 2007a). This framework has informed a variety of approaches to structuring online discussions, with the goal of enhancing the quality and inclusiveness of these spaces. Researchers have explored multiple avenues for improving discussion forums, including the development of structured workflows (Farnham et al., 2000; Kim et al., 2021a; Lee et al., 2020), the integration of  structural redesigns that facilitate more nuanced conversation (Zhang et al., 2017b; Tian et al., 2021; Nam and Ackerman, 2007; Rambow et al., 2004; Zhang and Cranshaw, 2018), and features that specifically facilitate the promotion of diverse viewpoints (Faridani et al., 2010; Munson et al., 2013; Kriplean et al., 2012a; Kim et al., 2021b). These efforts aim to create more organized, engaging, and meaningful discussions by combining the strengths of community-driven processes and automated tools.
@@
Previous studies have introduced various structural design changes aimed at improving navigation of and participation in long-tailed discussions. These structural representations included clustering, summaries, and threads to easily navigate and gain overview of the discussion. Clusters and threads have been found to be useful to identify insightful comments while navigating complex structures of discussion at varying levels of detail. For example, ordering comments into visual clusters (Hoque and Carenini, 2015) and hierarchical thematic organization (Hoque and Carenini, 2016) supported users in identifying insightful contributions and easily narrowing down to a subset of conversations. Adding a thread structure in a discussion space has been shown to increased user retention in comment participation (Budak et al., 2017), showing how structural design changes can also impact the user behavior. Summaries have played a crucial role in the synthesis of ideas within the discussion to encourage participants to reflect on the conversation (Kriplean et al., 2012b) and provide a structured overview that help readers navigate the main topics (Zhang et al., 2017b). Systems like Wikum (Zhang et al., 2017b) exemplify this approach by employing recursive summarization, enabling users to distill key points from extensive discussions. This method expanded to the creation of “living summaries” (Tian et al., 2021) that evolve as new contributions and insights emerge, ensuring that the discourse remains coherent and accessible. Beyond these approaches, some efforts have focused on developing lightweight tools to add structure that support easier contribution in the unstructured discussions spaces (Schneider et al., 2011; Zhang and Cranshaw, 2018). While these techniques have improved the accessibility of information by restructuring the discussion, their potential to actively foster the collaborative building of shared understanding and collective insights during the discussion process remains underexplored.
@@
Beyond structuring workflows, another crucial aspect of improving discussion spaces is promoting the inclusion of diverse viewpoints. Encouraging exposure to a variety of perspectives not only enriches the conversation but also helps users recognize and challenge their biases. Tools like Balancer (Munson et al., 2013) and ConsiderIt (Kriplean et al., 2012a) are designed to nudge users toward engaging with content from different sources and perspectives, fostering a more balanced and informed discourse. By actively recommending opposing viewpoints, systems like those developed by Gao et al. (Gao et al., 2018) and Nelimarkka et al. (Nelimarkka et al., 2019) mitigate selective exposure, encouraging users to explore opinions that differ from their own.  Systems such as Reflect (Kriplean et al., 2012b) promote active listening by summarizing others’ comments, leading to a deeper understanding of the intention of the commenter. In addition, interacting with moral framing grounded in frameworks such as Moral Foundations Theory (MFT) enables users to reflect on shared moral values underlying different perspectives, encouraging to rethink and engage with opposing viewpoints (Wang et al., 2022). Various visualization methods have also been used to map the diversity of users’ opinions and help navigate them effectively (Faridani et al., 2010; Kim et al., 2021b).
@@ 3.1. Design Inspiration from ‘SlashDot’
Our system is inspired by Slashdot, [^1] https://slashdot.org/ a platform known for its distributed approach to managing user behaviors in large-scale online discussions. Slashdot’s system allows users to take on specific roles in moderating the comments, with three levels of user participation: moderators, meta-moderators, and users. By decentralizing the moderation process to the user base, this approach has been successful in quickly and consistently distinguishing between high- and low-quality comments, reducing the burden on centralized staff to handle disruptive behaviors (Lampe and Resnick, 2004).
Slashdot’s approach of granting power to standout community members encouraged contributions that meet the community’s norms of quality (Kraut, 2012). By distributing responsibility among its members, it encouraged collective action to improve the discussion space. Inspired by this, our goal is to adapt a version of this distributed model to commenting systems in order to build a collective discussion space that incorporates the complexity of viewpoints, shifting the focus from fragmented individual expressions to collaboratively structured and managed discussions. This approach empowers users to take on diverse roles in organizing the discussion space, preventing dominance by any single viewpoint and ensuring discussions reflect a broad range of perspectives.
We propose a three-stage structure for collectively organizing discussions through distributed roles: clustering, summarization, and threading. Summaries have proven essential in previous research, synthesizing ideas with discussions and encouraging participants to reflect on the conversation (Zhang et al., 2017b; Nam and Ackerman, 2007), thus helping them actively engage with other’s perspectives (Kriplean et al., 2012b). By integrating these summaries with clustering and threading features, we transform flat discussions into a multi-level structure, allowing users to easily navigate between perspectives and explore ideas in greater depth (Hoque and Carenini, 2015, 2016). Clustering organizes related comments into groups, identifying recurring themes or shared viewpoints; summaries condense these clusters into concise insights; and threading organizes the clusters and summaries hierarchically. Through these features, the system collaboratively builds a collective understanding of complex viewpoints, creating a self-sustaining environment where every user contributes to the constructive exchange of opinions and maintains the quality of discussions through assigned activities, in line with a crowd-based moderation approach (Diakopoulos and Naaman, 2011).
@@ 3.2. Role Assignment
To implement a distributed model, we introduce three distinct user roles (Level 0, 1, and 2) that collaborate on structuring and organizing of discussions, focusing on three main features – clustering, summarizing, and threading. These features are not assigned individually to each level but are collectively realized through the collaboration of different levels. We distributed the roles so that the main outcomes of the system are realized through collaborative cycles of proposing and reviewing activities by users at different levels; for example, a Level0 user may propose a comment cluster, which is then reviewed by a Level1 user. The assigned roles for each level are presented in Figure 1.
Figure 1. Assigned roles for each user level: Lower levels contribute to smaller units of the discussion space, ordered as clusters, summarizations, and threads. All users can comment, while clustering, summarizing, and threading are collaboratively created and reviewed across different user levels. AI assists by suggesting summarizations and thread topics.:
@@
Level2 (LV2) users are responsible for creating new threads that introduce discussion topics and shape the high-level flow of the discussion. They can propose threads based on their perspectives, either branching off from existing discussions or addressing gaps in the current conversation, with AI suggestions assisting in the process. These proposed threads are then reviewed by other LV2 users, who decide whether to accept or decline them. The required number of approvals or denials for the creation of clusters and threads can be determined by considering the need for sufficient deliberation while avoiding significant delays in their creation. In this paper, we predefine this number as three considering the scale of the experimental environment.
@@ 3.4. System Features
Figure 2. The overview of our system: (a) guiding discussion topics, three of which are initially generated by AI, (b) discussion thread showing the overview of the corresponding comment section for each discussion topic, (c) summaries of clusters displayed in each discussion thread, along with timestamps for the summaries created, and (d) guiding question to prompt the discussion in each discussion thread, generated when the topic is created:
@@ 3.4.1. Clustering
To review these clustering activities, LV1 users can access the review page by clicking the “Review Clustered Comments” button at the top right of their screen. A list of cluster reviews will be displayed, allowing LV1 users to compare the discussion space before and after clustering. The left side shows the comments before clustering, while the right side displays the updated arrangement. During the review process, LV1 users can approve or decline clustering activities. A cluster is displayed in the comment section once it has been approved by the required number of LV1 users and is removed if it is declined by the predefined number of users.. In this paper, we set the requirement for both approval and denial to three participants, but the system can be adjusted to accommodate different thresholds based on their needs. The final clusters will be displayed in a blue box, visible to all users.
Figure 3. Workflow of Clustering: LV0 users propose clusters by dragging and dropping comments into the desired locations (left). LV1 users then review these clusters by comparing the changes before and after the clustering activity. The clusters become visible to all users once they are approved by the required number of LV1 users (right).:

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791518isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Collaborative and social computing systems and tools
@@ 1. Introduction
Previous studies have explored how the design of commenting systems can shape user behavior (Kiskola et al., 2021; Seering et al., 2019a). In many current news outlets, standardized commenting systems are used, where users simply post their comments or reply to others with limited opportunities to interact with other participants. This design often results in comments that are isolated expressions of personal thoughts, making it difficult for meaningful discussions to develop. As a result, discussions tend to unfold as a series of individual expressions, with users primarily focused on sharing their own views rather than collaborating toward a collective understanding or outcome. This behavior leads to fragmented conversations (Dahlberg, 2007a), where the overall discussion lacks depth. Consequently, similar viewpoints are often amplified, while diverse perspectives remain underrepresented (Anderson et al., 2014).
@@
Our research contributes to the exploration of an untapped space in designing collective online discourse by introducing a collaborative approach centered on “distributed roles,” reimagining how commenters can collectively shape and moderate discussions. We design and implement a novel system grounded in insights from previous works. Additionally, through experimentation, we present detailed findings on how our system differently shapes user behavior, understanding, and identifying the newly emerged behavioral patterns using our suggested system. Finally, through our design exploration, we identify key design considerations and trade-offs associated with the values embedded in our system, which can serve as guidance for future efforts in designing collective discourse systems.
@@ 2. Related Work
By inviting reader comments, today’s news media have embraced greater user involvement in the journalistic process. This phenomenon has been explored by scholars through frameworks such as “participatory journalism” (Singer and Reich, 2011) and “reciprocal journalism” (Lewis et al., 2014); the audience has also become an active contributor to content. While there were positive views of news commenting acting as the digital cafés of a Public Sphere 2.0 (Ruiz et al., 2011), journalists and news outlets were often concerned about quality control, manageability, and the maintenance costs of user-generated content (Diakopoulos and Naaman, 2011).
@@ 2.2. Improving Discussion Quality through Moderation
A pre-moderation approach seeks to improve the discussion quality by addressing potential issues before they arise. These proactive approaches range from using design elements to nudge users toward prosocial participation (Seering et al., 2019a), to more direct interventions that review and approve comments before they appear publicly. Common design elements include explicitly displaying community rules and guidelines (Kraut and Resnick, 2012; Matias, 2019a), using prompts to encourage more thoughtful participation (McInnis et al., 2016), and employing curation strategies that influence users’ deliberation (McInnis et al., 2018). More direct strategies involve using algorithmic support to provide risk information to users (Chang et al., 2022) or moderators (Schluger et al., 2022), helping in the identification and resolution of tensions before they escalate.
The post-moderation approach, by contrast, takes place after content has been published. It focuses on identifying and mitigating the impact of harmful or inappropriate material through reactive review of comments. While traditional methods involve human moderators reviewing flagged content (Roberts, 2019; Seering et al., 2019b), the scale and volume of content necessitates heavy reliance on community participation to flag or report inappropriate comments. Community-driven methods, also called a crowd-based approach, have proven to be effective for filtering and evaluating the quality of comments (Diakopoulos and Naaman, 2011). This approach leverages the collective judgment of the community to assess and organize comments, fostering a more dynamic and responsive environment. Collaborative platforms such as Wikipedia exemplify this, where community members work together to resolve disputes (Im et al., 2018) and contribute to the creation of high-quality articles (Zhang et al., 2017b). Another notable example is the moderation system used by Slashdot (Lampe et al., 2007), where users rate and moderate comments, with these decisions aggregated to determine comment quality. Slashdot’s effectiveness lies in its ability to harness the collective intelligence of its users, allowing for real-time participation in organizing and structuring discussions. This distributed form of moderation is both scalable and adaptable, addressing the diverse needs of its community.
@@ 2.3. Restructuring and Reimagining the Discussion Space
The deliberative model of democracy emphasizes the importance of reasoned argumentation, mutual respect, and the willingness to consider others’ perspectives as essential components for effective discourse (Dahlberg, 2007b). This framework has informed a variety of approaches to structuring online discussions, with the goal of enhancing the quality and inclusiveness of these spaces. Researchers have explored multiple avenues for improving discussion forums, including the development of structured workflows (Farnham et al., 2000; Kim et al., 2021b; Lee et al., 2020), the integration of structural redesigns that facilitate more nuanced conversation (Zhang et al., 2017a; Tian et al., 2021; Nam and Ackerman, 2007; Rambow et al., 2004; Zhang and Cranshaw, 2018), and features that specifically facilitate the promotion of diverse viewpoints (Faridani et al., 2010; Munson et al., 2013; Kriplean et al., 2012a; Kim et al., 2021a). These efforts aim to create more organized, engaging, and meaningful discussions by combining the strengths of community-driven processes and automated tools.
@@
Previous studies have introduced various structural design changes aimed at improving navigation of and participation in long-tailed discussions. These structural representations included clustering, summaries, and threads to easily navigate and gain overview of the discussion. Clusters and threads have been found to be useful to identify insightful comments while navigating complex structures of discussion at varying levels of detail. For example, ordering comments into visual clusters (Hoque and Carenini, 2015) and hierarchical thematic organization (Hoque and Carenini, 2016) supported users in identifying insightful contributions and easily narrowing down to a subset of conversations. Adding a thread structure in a discussion space has been shown to increased user retention in comment participation (Budak et al., 2017), showing how structural design changes can also impact the user behavior. Summaries have played a crucial role in the synthesis of ideas within the discussion to encourage participants to reflect on the conversation (Kriplean et al., 2012b) and provide a structured overview that help readers navigate the main topics (Zhang et al., 2017a). Systems like Wikum (Zhang et al., 2017a) exemplify this approach by employing recursive summarization, enabling users to distill key points from extensive discussions. This method expanded to the creation of “living summaries” (Tian et al., 2021) that evolve as new contributions and insights emerge, ensuring that the discourse remains coherent and accessible. Beyond these approaches, some efforts have focused on developing lightweight tools to add structure that support easier contribution in the unstructured discussions spaces (Schneider et al., 2011; Zhang and Cranshaw, 2018). While these techniques have improved the accessibility of information by restructuring the discussion, their potential to actively foster the collaborative building of shared understanding and collective insights during the discussion process remains underexplored.
@@
Beyond structuring workflows, another crucial aspect of improving discussion spaces is promoting the inclusion of diverse viewpoints. Encouraging exposure to a variety of perspectives not only enriches the conversation but also helps users recognize and challenge their biases. Tools like Balancer (Munson et al., 2013) and ConsiderIt (Kriplean et al., 2012a) are designed to nudge users toward engaging with content from different sources and perspectives, fostering a more balanced and informed discourse. By actively recommending opposing viewpoints, systems like those developed by Gao et al. (Gao et al., 2018) and Nelimarkka et al. (Nelimarkka et al., 2019) mitigate selective exposure, encouraging users to explore opinions that differ from their own. Systems such as Reflect (Kriplean et al., 2012b) promote active listening by summarizing others’ comments, leading to a deeper understanding of the intention of the commenter. In addition, interacting with moral framing grounded in frameworks such as Moral Foundations Theory (MFT) enables users to reflect on shared moral values underlying different perspectives, encouraging to rethink and engage with opposing viewpoints (Wang et al., 2022). Various visualization methods have also been used to map the diversity of users’ opinions and help navigate them effectively (Faridani et al., 2010; Kim et al., 2021a).
@@ 3.1. Design Inspiration from ‘SlashDot’
Our system is inspired by Slashdot [^1] https://slashdot.org/, a platform known for its distributed approach to managing user behaviors in large-scale online discussions. Slashdot’s system allows users to take on specific roles in moderating the comments, with three levels of user participation: moderators, meta-moderators, and users. By decentralizing the moderation process to the user base, this approach has been successful in quickly and consistently distinguishing between high- and low-quality comments, reducing the burden on centralized staff to handle disruptive behaviors (Lampe and Resnick, 2004).
Slashdot’s approach of granting power to standout community members encouraged contributions that meet the community’s norms of quality (Kraut and Resnick, 2012). By distributing responsibility among its members, it encouraged collective action to improve the discussion space. Inspired by this, our goal is to adapt a version of this distributed model to commenting systems in order to build a collective discussion space that incorporates the complexity of viewpoints, shifting the focus from fragmented individual expressions to collaboratively structured and managed discussions. This approach empowers users to take on diverse roles in organizing the discussion space, preventing dominance by any single viewpoint and ensuring discussions reflect a broad range of perspectives.
We propose a three-stage structure for collectively organizing discussions through distributed roles: clustering, summarization, and threading. Summaries have proven essential in previous research, synthesizing ideas with discussions and encouraging participants to reflect on the conversation (Zhang et al., 2017a; Nam and Ackerman, 2007), thus helping them actively engage with others’ perspectives (Kriplean et al., 2012b). By integrating these summaries with clustering and threading features, we transform flat discussions into a multi-level structure, allowing users to easily navigate between perspectives and explore ideas in greater depth (Hoque and Carenini, 2015, 2016). Clustering organizes related comments into groups, identifying recurring themes or shared viewpoints; summaries condense these clusters into concise insights; and threading organizes the clusters and summaries hierarchically. Through these features, the system collaboratively builds a collective understanding of complex viewpoints, creating a self-sustaining environment where every user contributes to the constructive exchange of opinions and maintains the quality of discussions through assigned activities, in line with a crowd-based moderation approach (Diakopoulos and Naaman, 2011).
@@ 3.2. Role Assignment
To implement a distributed model, we introduce three distinct user roles (Level 0, 1, and 2) that collaborate on structuring and organizing of discussions, focusing on three main features – clustering, summarizing, and threading. These features are not assigned individually to each level but are collectively realized through the collaboration of different levels. The role hierarchy in our system was designed so that users at each level take responsibility for increasingly broader units of the discussion. Level 0 users start by clustering individual comments. Level 1 users then review and summarize each cluster, working with ideas at a wider scope. Finally, Level 2 users create and organize threaded discussions based on those summaries. Assigning roles in this order ensures that each level’s participation is enabled by the validated outputs from the previous level and that collaboration emerges through cycles of proposing and reviewing across role boundaries.
The assigned roles for each level are presented in Figure 1.
Figure 1. Assigned roles for each user level: Lower levels contribute to smaller units of the discussion space, ordered as clusters, summarizations, and threads. All users can comment, while clustering, summarizing, and threading are collaboratively created and reviewed across different user levels. AI assists by suggesting summarizations and thread topics.
The figure shows a diagram of user roles divided by level, including Level 0, Level 1, and Level 2, where each level has specific actions related to commenting, clustering, summarizing, and threading. Lower levels contribute to smaller units of the discussion space, ordered as clusters, summarizations, and threads. At the top, all levels are shown able to ’Write’ in the comment section. Beneath that, to make a cluster, Level 0 users can ’Create’ and Level 1 users can ’Review.’ To create a summarization, Level 1 users can ’Create’ with AI suggesting the summary in this process. To make a new thread, Level 2 users can ’Create’ the thread topic and ’Review’ the created topics, with AI also suggesting topics in this process.
@@
Level2 (LV2) users are responsible for creating new threads that introduce discussion topics and shape the high-level flow of the discussion. They can propose threads based on their perspectives, either branching off from existing discussions or addressing gaps in the current conversation, with AI suggestions assisting in the process. These proposed threads are then reviewed by other LV2 users, who decide whether to accept or decline them. The required number of approvals or denials for the creation of clusters and threads can be determined by considering the need for sufficient deliberation while avoiding significant delays in their creation. In this paper, we predefine this number as three, considering the scale of the experimental environment.
@@ 3.4. System Features
Figure 2. The overview of our system: (a) guiding discussion topics, three of which are initially generated by AI, (b) discussion thread showing the overview of the corresponding comment section for each discussion topic, (c) summaries of clusters displayed in each discussion thread, along with timestamps for the summaries created, and (d) guiding question to prompt the discussion in each discussion thread, generated when the topic is created
The figure shows an interface for the system overview. At the top, three (a) guiding topics are displayed: ’Olympic Games Housing Impact’, ’Homelessness and Urban Development’, and ’Social Responsibility in Olympics’, which are initially generated by AI. Beneath the guiding topics, there are corresponding (b) discussion threads, each showing the overview of a corresponding comment section. In these discussion threads, (c) summaries of clustered comments are shown along with the timestamps of their creation. In the figure, ’Balancing social responsibility with business interest at Olympic’ is highlighted as an example of summary. After clicking on one of the discussion threads (e.g., ’Homelessness and Urban Development), the comment section appears, with (d) guiding questions at the top (e.g., ’What ethical considerations should be made when urban development displaces homeless populations?’), which are designed to prompt the discussion and are initially generated when the topics are created.
@@ 3.4.1. Clustering
To review these clustering activities, LV1 users can access the review page by clicking the “Review Clustered Comments” button at the top right of their screen. A list of cluster reviews will be displayed, allowing LV1 users to compare the discussion space before and after clustering. The left side shows the comments before clustering, while the right side displays the updated arrangement. During the review process, LV1 users can approve or decline clustering activities. A cluster is displayed in the comment section once it has been approved by the required number of LV1 users and is removed if it is declined by the predefined number of users. In this paper, we set the requirement for both approval and denial to three participants, but the system can be adjusted to accommodate different thresholds based on their needs. The final clusters will be displayed in a blue box, visible to all users.
Figure 3. Workflow of Clustering: LV0 users propose clusters by dragging and dropping comments into the desired locations (left). LV1 users then review these clusters by comparing the changes before and after the clustering activity. The clusters become visible to all users once they are approved by the required number of LV1 users (right).
The figure shows the workflow of clustering in a system where Level 0 (LV0) and Level 1 (LV1) users collaborate together to create a cluster. The left side of the figure illustrates how LV0 proposes a cluster by dragging and dropping comments into desired locations. The interface displays individual comments, with LV0 users arranging them into clusters, which remain visible only to the creator during this stage. On the right side, after the clustering activity, LV1 users review the clusters by comparing the changes before and after clustering, with a side-by-side view of comments showing the ’Before’ and ’After’ state of clustering. LV1 users can either approve or decline each cluster, and once a certain number (N=3) of LV1 users agree, the clusters become visible to all users across levels (LV0, LV1, LV2). The figure includes step-by-step instructions for this process: For LV0 users, (1) drag and drop the comments, and (2) the cluster remains visible only to the creator. For LV1 users, (1) click the ’Review Clustered Comments’ button, (2) review the clusters by clicking ’Approve’ or ’Decline’ button, and (3) the clusters become visible to other levels after approved by the required number of LV1 users.